OpenStudy (anonymous):

Find the inverse of the function, f ^−1(x), if it exists. f(x)=-2-5x/4+5x assume that x DOES NOT = -0.8

OpenStudy (anonymous):

x=-2-5y/4+5y Isolate the variable's so x+2=20/4y-5/4y x+2=15/4y divide by 15/4 is the same as multiply by 4/15 so you are left with y=4x+8/15